OpenStudy (anonymous):

solve 7(x+2)-10=7x+4

OpenStudy (anonymous):

7(x+2)-10=7x=4 Is this Distibutive Property?

OpenStudy (horsegirl27):

I can help

OpenStudy (horsegirl27):

yes distribute 7

OpenStudy (horsegirl27):

Or I can just explain it quickly

OpenStudy (horsegirl27):

Either one

OpenStudy (anonymous):

Let @horsegirl27 explain it

OpenStudy (horsegirl27):

7(x+2)-10=7x+4 7x+14-10=7x+4

OpenStudy (anonymous):

7(x+2)-10=7x+4 7x +14-10=7x+4 and i think u subrtact 7 14-10=4 4=4

OpenStudy (horsegirl27):

Ok those are the first steps

OpenStudy (anonymous):

Is that rght?

OpenStudy (horsegirl27):

Yes I think so

OpenStudy (anonymous):

there is no solution i just tried it for myself and there is nothing

OpenStudy (horsegirl27):

@cashmoney3456 what are the options? Is it multiple choice?

OpenStudy (anonymous):

there is no solution because the x cancels out

OpenStudy (horsegirl27):

yes

OpenStudy (anonymous):

u got it